# My CXL Pool Obviates Your PCIe Switch

### v2

@@ Abstract.
ccs: Software and its engineering Operating systemsccs: Computer systems organization Cloud computingccs: Hardware Emerging technologies
@@ 1. Introduction
  • Utilization. Pooling PCIe devices mitigates stranded I/O resources and improve device utilization, which reduces datacenter TCO (pond, ; coach, ).
@@
In this paper, we argue that PCIe device pooling can be implemented in software on top of Compute eXpress Link (CXL) memory pools. CXL memory pools have gained industry interest to improve memory utilization and scale up in-memory databases (ahn2024examination, ; pond, ; hyperscale2023cxl, ; seagate2024cxl, ; ha2023dynamic, ; marvell2024structera, ; ufx-news, ; astera-news, ; cxl-db, ; lerner2024cxl, ). A CXL pod (witchel2024challenges, ) consists of multiple hosts within a rack, allowing these hosts to dynamically allocate memory from the pool based on demands. Recent work shows how to build CXL pods with hardware available today for about \$$\$600$ per host (berger2025octopus, ). This enables positive return-on-investment (ROI) for memory-pooling use cases (berger2023design, ).
@@
In this paper, we first motivate PCIe device pooling (§2) and provide background on CXL memory pools (§3). We then sketch our design of software-based PCIe pooling with CXL (§4). Lastly, we discuss how PCIe device pooling will change the datacenter infrastructure and other open questions (§5).
@@ 3. CXL Memory Pools
CXL memory pools can be constructed either by CXL devices with multiple ports or by CXL switches. Devices with multiple CXL ports, or multi-headed devices (MHD), can connect to multiple hosts via dedicated CXL links (pond, ; ufx-news, ; astera-news, ; berger2025octopus, ; hyperscale2023cxl, ; marvell2024structera, ; ha2023dynamic, ; seagate2024cxl, ; mackey2024cxl, ). A single MHD today has up to 20 CXL ports (ufx-news, ) and multiple MHDs can be combined to scale to larger CXL pod sizes (berger2025octopus, ). CXL switches can scale to significantly larger pod sizes (xconn-cxl, ), e.g., 4096 in CXL 3.0 (cxl-intro, ; cxl-3.0, ). However, CXL switches are more costly and MHD-based pods will likely be deployed earlier (berger2023design, ; against-cxl, ; berger2025octopus, ).
@@
A CXL memory pool can be used as shared memory across all the hosts in a CXL pod. Shared CXL memory has the potential to speed up RPCs, key-value stores, and distributed databases (hydrarpc, ; rpcool, ; ahn2024examination, ; cxl-shm, ),
@@ 4.1. PCIe Datapath Over CXL
For non-cache-coherent CXL pools, the datapath should explicitly maintain coherency in software. When modifying I/O buffers, the data should always be written to the CXL memory rather than staying in the CPU caches. Otherwise, other hosts might retrieve stale data from the CXL memory.
@@
To quantify if placing I/O buffers in CXL affects I/O latency, we set up a two-socket server equipped with a 100 Gbps Mellanox ConnectX-5 NIC and connect both of its CPUs to a MHD-based CXL pod with each a PCIe-5.0 $\times$8 link. A second host serves as load generator (client) and connects via another 100 Gbps ConnectX-5 NIC to a common 100 Gbps switch. We use Junction (junction, ) as the network stack for both the server and the client. We modify Junction on the server side to allocate TX and RX buffers from the CXL memory pool. The NIC connects to socket0 and uses one $\times$8 CXL link. Junction runs on socket1 and uses the other $\times$8 CXL link.
@@ 4.2. Pooling Orchestrator
Each host runs a pooling agent that monitors and configures the PCIe device. The orchestrator and the agents communicate using shared-memory channels in the shared CXL memory. Both roles are good candidate to offload to accelerators such as SmartNICs.
@@ 5. Discussion
Such an approach could also extend to storage devices. For example, a storage server in an object storage service like S3 could shift load across a large number of SSDs if it is writing a large amount of data requiring high storage bandwidth. This may behave like adaptive storage striping or RAID configurations.

### v3

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: Workshop in Hot Topics in Operating Systems; May 14–16, 2025; Banff, AB, Canadabooktitle: Workshop in Hot Topics in Operating Systems (HOTOS 25), May 14–16, 2025, Banff, AB, Canadadoi: 10.1145/3713082.3730393isbn: 979-8-4007-1475-7/25/05ccs: Software and its engineering Operating systemsccs: Computer systems organization Cloud computingccs: Hardware Emerging technologies
@@ 1. Introduction
  • Utilization. Pooling PCIe devices mitigates stranded I/O resources and improve device utilization, which reduces datacenter total cost of ownership (TCO) (pond, ; coach, ).
@@
In this paper, we argue that PCIe device pooling can be implemented in software on top of Compute eXpress Link (CXL) memory pools. CXL memory pools have gained industry interest to improve memory utilization and scale up in-memory databases (ahn2024examination, ; pond, ; hyperscale2023cxl, ; seagate2024cxl, ; ha2023dynamic, ; marvell2024structera, ; ufx-news, ; astera-news, ; cxl-db, ; lerner2024cxl, ). Hardware vendors have also started offering CXL memory with pooling capabilities [^2] Although advanced pooling features (e.g., Dynamic Capacity Device (DCD), Back Invalidation (BI)) are only standardized in the recent CXL 2.0 and 3.0 specification (cxl-3.0, ; cxl-2.0, ), the widely adopted CXL 1.1 specification already provides the necessary functionalities for CXL memory pools (§3) (ufx-news, ; astera-news, ; marvell2024structera, ). A CXL pod (witchel2024challenges, ) consists of multiple hosts within a rack, allowing these hosts to dynamically allocate memory from the pool based on demands. Recent work shows how to build CXL pods with hardware available today for about \$$\$600$ per host (berger2025octopus, ) without using expensive CXL switches (§3). This enables positive return-on-investment (ROI) for memory-pooling use cases (berger2023design, ).
@@
In this paper, we first motivate PCIe device pooling (§2) and provide background on CXL memory pools (§3). We then sketch our design of software-based PCIe pooling with CXL (§4). Lastly, we discuss how PCIe device pooling will change the datacenter infrastructure and other open questions, including how the CXL link bandwidth affects different PCIe pooling use cases (§5).
@@ 3. CXL Memory Pools
CXL memory pools can be constructed either by CXL devices with multiple ports or by CXL switches. Devices with multiple CXL ports, or multi-headed devices (MHD), can connect to multiple hosts via dedicated CXL links (pond, ; ufx-news, ; astera-news, ; berger2025octopus, ; hyperscale2023cxl, ; marvell2024structera, ; ha2023dynamic, ; seagate2024cxl, ; mackey2024cxl, ). These devices are readily available today from AsteraLabs (leo2023, ), Marvell (marvell2024structera, ), and UnifabriX (ufx-news, ). A single MHD today has up to 20 CXL ports (ufx-news, ) and multiple MHDs can be combined to scale to larger CXL pod sizes (berger2025octopus, ).
On the other hand, CXL switches allow CXL devices (including devices with a single port) to be connected to a CXL switch, which in turn connects to multiple hosts. Switched designs resemble tree network topologies in CXL 2.0 (cxl2topology, ) and Clos network topologies in CXL 3.0 (cxl3fabric, ). A single switch typically offers 128-256 CXL lanes (xconn-cxl, ). These lanes would be shared among connected CPUs, CXL controllers, and other devices. For example, 128 lanes could connect six CPUs with x16 CXL ports and two single-ported CXL controllers with x16 ports. In CXL 3.0, multiple switching levels can be used to build large-scale CXL pods with 4096 hosts (cxl-intro, ).
The tradeoff for switched designs is higher power, cost, and latency. Prior work observes that CXL switch costs can exceed the benefits that can be achieved from pooling at scale (berger2023design, ; against-cxl, ). CXL switch latency is also fundamentally higher than direct CXL connections, as switches require serialization and deserialization of CXL packets twice on every path from CPU to CXL memory controller (berger2023design, ; pond, ). Current switches add more than $250ns$ of latency (xconn2024latency, ), resulting in idle load-to-use latency of roughly $500-600ns$.
In short, CXL switches are more costly and have higher latency overhead, so MHD-based pods will likely be deployed earlier (berger2023design, ; against-cxl, ; berger2025octopus, ).
@@
A CXL memory pool can be used as shared memory across all the hosts in a CXL pod. Shared CXL memory has the potential to speed up RPCs, key-value stores, and distributed databases (hydrarpc, ; rpcool, ; ahn2024examination, ; cxl-shm, ; tigon, ; pasha, ; ctxnl, ; cxl-sync, ),
@@ 4.1. PCIe Datapath Over CXL
For non-cache-coherent CXL pools, the datapath should explicitly maintain coherence in software. When modifying I/O buffers, the data should always be written to the CXL memory rather than staying in the CPU caches. Otherwise, other hosts might retrieve stale data from the CXL memory.
@@
To quantify if placing I/O buffers in CXL affects I/O latency, we set up a two-socket server equipped with a 100 Gbps Mellanox ConnectX-5 NIC and connect both of its CPUs to a MHD-based CXL pod with each a PCIe-5.0 $\times$8 link. A second host serves as load generator (client) and connects via another 100 Gbps ConnectX-5 NIC to a common 100 Gbps switch. We use Junction (junction, ) as the network stack for both the server and the client. We modify Junction on the server side to allocate TX and RX buffers (not the TX/RX queues) from the CXL memory pool. The NIC connects to socket0 and uses one $\times$8 CXL link. Junction runs on socket1 and uses the other $\times$8 CXL link.
@@ 4.2. Pooling Orchestrator
Each host runs a pooling agent that monitors and configures the PCIe device. The orchestrator and the agents communicate using shared-memory channels in the shared CXL memory. Both roles are good candidates to offload to accelerators such as SmartNICs.
@@ 5. Discussion
CXL link bandwidth
Recently platforms (e.g., Intel Xeon 6) provide 64 CXL 2.0 / PCIe5 lanes per CPU socket (lenovo2024thinksystem, ; asrock2024gnrd8, ; kalodrich2024supermicro, ; kaytus2024kr2280v3, ; kennedy2024lenovo, ). Fully disaggregating a 200 Gbps NIC and a 400 Gbps NIC requires only 8 and 16 CXL lanes, respectively. Therefore, pooling NICs to improve NIC utilization and reduce the NIC-to-host ratio can be supported by allocating enough CXL lanes per host to interleave between multiple MHDs. Similarly, disaggregating six NVMe SSDs (current AWS servers offer six local NVMe SSDs (awsnetwork, ), and datacenter SSDs today often provide 5 GB/s bandwidth (solidigmd5, ; 990-EVO, )) would require 30 GB/s bandwidth, which can be satisfied with 8 CXL lanes.
However, enabling a single host to saturate a large number of NICs to maximize its peak network performance does pose higher requirements for the CXL link (dfabric, ). For example, to drive the combined bandwidth of eight 400 Gbps NICs in a CXL pod, the host would need at least 100 CXL 2.0 lanes, making this use case less realistic.
Operational implications
Datacenters often need to update host OSes and platform firmware, which requires rebooting individual hosts. To enable such maintenance, PCIe device pooling should support hot-adding and hot-removing hosts from a CXL pod, allowing updates to be rolled out incrementally on a per-host basis. For instance, when reconfiguring a host in public clouds, the VM scheduler should live migrate all VMs off the host and notify the pooling orchestrator to hot-remove it. Upon removal, the orchestrator should prevent new device allocations to the host and migrate its existing PCIe device assignments to other active hosts.
6. Acknowledgments
We thank the reviewers for their helpful comments. We also thank Rodrigo Fonseca and Stefan Saroiu for their feedback. This work was supported by NSF award CNS-2143868.